Circular No. 7691

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dw IN NGC 1168 G. M. Hurst, Basingstoke, England, reports the discovery of an appar-

ent supernova (mag 16.6) by T. Boles, Coddenham, England, on unfiltered CCD images obtained on Aug. 25.125 (limiting mag 19) and 25.989 UT (poor conditions) with a 0.36-m Schmidt-Cassegrain telescope for the U.K. Nova/Supernova Patrol. The new object is located at α = 3^{h}00^{m}47^{s}.88, δ = +11^{o}46′27′′.5 (equinox 2000.0), which is 13′′ west and 9′′.5 south of the center of NGC 1168. A confirming image obtained by A. Oksanen, Nyrola Observatory, Finland, on Aug. 26.025 with a 0.36-m Schmidt-Cassegrain telescope yields the following position end figures for SN 2001dw: 47^{s}.84, 26′′.7. The new object was absent from a master image taken by Boles on 1999 Nov. 2, and it does not appear on second Palomar Sky Survey images from 1990 (blue) and 1991 (red).

V2275 CYGNI = NOVA CYGNI 2001 No. 2 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V2275 Cyg has been given to the nova reported on IAUC 7686, 7687, and 7688.

T. Kato, Kyoto University, reports the following additional magnitudes for V2275 Cyg: Aug. 17.615 UT, [11.5 (H. Yoshida, Sapporo, Hokkaido, Japan; 55-mm f/2.8 camera lens + Fujichrome TREBI 400 film); 17.585, [8.6: (Yoshida; emulsion uncertain); 18.506, 8.8 (Yoshida; 35-mm f/2.8 camera lens + Kodak Gold-Max 800 film); 18.512, 9.0 (H. Nishimura, Kakegawa, Shizuoka, Japan; 105-mm camera lens + T-Max 400 film); 18.544, 8.8 (Yoshida; Gold-Max 800 film).

Additional CCD V magnitudes (S = G. Sostero and A. Lepardo, Remanzacco, Italy; R = D. Rodriguez, Madrid, Spain; W = J. D. West, Mulvane, KS): Aug. 20.83 UT, 7.36 (S); 20.875, 6.7 (R); 21.090, 7.66 (W); 21.82, 8.05 (S); 22.106, 8.02 (W); 23.82, 8.55 (S); 25.01, 9.07 (S); 25.106, 9.01 (W; clouds); 26.01, 9.26 (S). Cousins I magnitudes by West: Aug. 21.088, 5.53; 22.107, 5.83; 25.094, 7.00 (clouds). Color photometry by Sostero and Lepardo: Aug. 20.83, B − V = +1.05, V − R_{c} = +0.7, R_{c} − I_{c} = +1.2; 21.82, +1.04, +0.91, +1.24; 23.82, +1.0, +1.0, +1.1; 25.01, +1.0, +1.1, +1.0; 26.01, +0.95, +1.1, +1.0.

2001 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green